# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ 6.5.  Finishing the proof of the full characterisation
We first check that if $m,n$ are of the shape of one of the cases (i)(iv), then $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective. In the cases (i), (ii) we have $m<q_{M+1}$, and so Lemma 6.5 implies that $n\in\{x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}\}$, and Lemma 6.4 implies that the corresponding function $f$ is bijective. If we are in the case (iii) and $n^{[\leq M-1]}=0$, then again Lemmas 6.5 and 6.4 imply that $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective. If we are in the case (iii) $n^{[\leq M-1]}>0$, this is provided by Lemma 6.12. If we are in the case (iv) and not in one of the previous cases, then Lemma 6.13 says that $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective.
For the implication in the other direction, assume that $2\leq m\leq n$ and that $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective. If $n\in\{x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}\}$, then Lemmas 6.4 and 6.5 (and the property (5.6)) imply that $m,n$ are of the shape (i) or (ii). If $n\neq x_{\textnormal{left}}^{*},x_{\textnormal{right}}^{*}$, then Lemma 6.11 implies that $m$ must either be a convergent or a semi convergent. If $m$ is a convergent, then Lemma 6.12 implies that $m,n$ are of the shape (iii) (or possibly (i)). If $m$ is a semi-convergent, then Lemma 6.13 implies that $m,n$ are of the shape (iv) (or possibly (i) or (ii)). ∎

### v2

@@ 6.5.  Very small $m$
We first check that if $m,n$ are of the shape of one of the cases (i)(iv), then $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective. In the cases (i), (ii) we have $m<q_{M+1}$, and so Lemma 6.5 implies that $n\in\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$, and Lemma 6.4 implies that the corresponding function $f_{*}$ is bijective. If we are in the case (iii) and $n^{[\leq M-1]}=0$, then again Lemmas 6.5 and 6.4 imply that $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective. If we are in the case (iii) and $n^{[\leq M-1]}>0$, this is provided by Lemma 6.13. If we are in the case (iv) and not in one of the previous cases, then Lemma 6.14 says that $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective.
For the implication in the other direction, assume that $2\leq m\leq n$ and that $f_{\textnormal{left}},f_{\textnormal{right}}$ are bijective.
If $n\in\{\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}\}$, then Lemma 6.5 and the property (5.6) imply that $m,n$ are of the shape (i) or (ii), or $m=q_{M}$. If $m$ is a convergent, then Lemma 6.13 implies that $m,n$ are of the shape (iii) or (i).
If $n\neq\widehat{x}_{\textnormal{left}},\widehat{x}_{\textnormal{right}}$, then Lemma 6.12 implies that $m$ must either be a convergent or a semi convergent, or $m<q_{1}$. If $m$ is a convergent, then as before Lemma 6.13 implies that $m,n$ are of the shape (iii) or (i). If $m$ is a semi-convergent, then Lemma 6.14 implies that $m,n$ are of the shape (iv) or (i) or (ii). Finally, if $m<q_{1}$, Lemma 6.15 implies that $m,n$ are of the shape (i) or (ii). ∎